OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2? y = 2x + 1 y = 2x + 3 y = 2x - 8 y = 2x - 3

OpenStudy (anonymous):

I think 2...not sure.

OpenStudy (anonymous):

Thanks for trying!

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

y = 2x + 1 y -5 = 2 ( x -2) ->y = 2x + 1